OpenStudy (anonymous):

I need some help on finding the solution, an ordered pair using the substitution method 3m + n = 1 m-5n = 11 can anyone help and explain I get so far and I am confused

OpenStudy (anonymous):

The substitution method means you solve one equation for a variable and then plug that solution into the other equation for that variable.

OpenStudy (anonymous):

so n = 1-3m m-5(1-3m) = 11 m-5+15m = 11 16m = 16 m=1. n = 1-3(1) = -2 n=-2

OpenStudy (anonymous):

this is where i get confused

OpenStudy (anonymous):

What are you confused on? Once you have your answers i.e m=1 and n=-2 you can write them as an ordered pair.

OpenStudy (anonymous):

3m + n = 1 subtract 3m from both sides n=1-3m what r u confuzed about?

OpenStudy (anonymous):

So (n,m) could be a pair or (m,n) depends on which one is independent and dependent.

OpenStudy (anonymous):

why do you pick 3m, is there a reason on hwich one to pick

OpenStudy (anonymous):

He is just solving for one of the variables if that is what you are asking. You want to get one of the variables in terms of the other so you will have an equation with only one unknown that you can solve for. He could have done 3m=1-n and then divided by 3 to solve for m and plug in into the other equation.

OpenStudy (anonymous):

3m + n = 1 m - 5n = 11 3m - 3m + n = 1 - 3m

OpenStudy (anonymous):

is that what i suppossed to do, then that equation would become n = 1-3m

OpenStudy (anonymous):

That is how you would solve the first equation for n. Now you would replace n in the other equation with your "new" n=1-3m so m-5(1-3m)=11

OpenStudy (anonymous):

m - 5(1-3m) = 11 this correct? then I would make it 5 - 16m = 11

OpenStudy (anonymous):

The first part is correct, but the second part I believe you messed up on your signs. Should be -5+16m=11

OpenStudy (anonymous):

Then how do I go next

OpenStudy (anonymous):

Add 5 to both sides to get 16m=16 divide both sides by 16 to leave m=1

OpenStudy (anonymous):

oh I thought I would have to subtract 5 from both sides not add, this is where I get confused on knowing which to do , add or subtract

OpenStudy (anonymous):

Since the 5 is negative, you want to perform the opposite operation to get it on the other side. If the 5 was positive you would subtract it from both sides.

OpenStudy (anonymous):

thank you this helps very much

OpenStudy (anonymous):

You're welcome!